DIRC Operations Weekly Report

Period: Saturday 2006/06/23 - Friday 2006/06/30

Oncall: Rolf Andreassen, Jerome Bequilleux
Fast Monitoring: Rolf Andreassen, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060630.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060630 (empty)

Contents


BaBar-PEP news

PEP peak lumi: xxx*1033 (previous week: 10.2*1033)

BaBar integrated lumi: xxxx fb-1 (previous week: 2.831 fb-1)

Run 5b PEP delivery: 62.49 fb-1

Run 5 total delivery: 120.72 fb-1 Run 5 total BaBar recorded: 116.35 fb-1 including 9.82 fb-1 off-peak data Note : No PEP-BaBar meeting last week; numbers in red are estimated 'by hand' from plots. Numbers in black are official.

PEP activities this week:

  • UPS replacement on Monday led to about five hours downtime, and was not successful.

    PEP activities planned next week :

  • Foreseen work opportunities (PEP down periods):

  • BaBar activities/issues this week:

  • DIRC activities

    See DIRC logbook for details
  • Friday 23rd : Matt replaced a water pump. Discovered that fan monitoring plots were not being updated properly; it seems the commissioners have failed to look at the dates on the x axis. The 'monthly' plots were showing April... Nicolas fixed the scripts. Lesson : Do not just look at the time development, also check what time the plots are for!
  • Monday 26th : To avoid problems while the UPS work was going on, Matt disabled the dump valve and put the degassing and vacuum pumps on house power. When the work ended, we found that the VME crate was not communicating with epics; we replaced the fan tray, and that fixed thsi problem. We also needed to reboot the GMB.
  • Wednesday 28th : Jerome takes over oncall, Rolf goes off on vacation.

    Downtime due to the DIRC: None.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Links